OpenStudy (anonymous):

how to find the exact value of 23pi/6

Parth (parthkohli):

Using a calculator.

OpenStudy (anonymous):

what do i do if it says not to use one?

OpenStudy (anonymous):

Try to approximate it as closely as possible. Familiarize yourself with the unit circle.

OpenStudy (anonymous):

\[\frac{ 23\pi }{ 6 }=\frac{ 23(180) }{ 6 }\]

OpenStudy (anonymous):

\[\pi=180\]

Parth (parthkohli):

\[\pi = 180^{\circ}\]Yes.

OpenStudy (anonymous):

I thought this guy wanted to know sin/cos/tan(23pi/6). But now I think he meant to find the circumference or something.